# Federated Neural Graph Databases

### v3

@@ Appendix A Parameter Sharing
In this section, we provide an example of sharing secrets between clients. The parameters can be shared under the protection using various encryption methods, for example, the commonly used is Diffie–Hellman key exchange (Diffie and Hellman, 2022) shown as follows, we consider the sharing process between two clients:
@@ 2  Transmit $\theta_{i}^{r}$ to all other clients with encryption, each
B.2. FedNGDB Framework
We present the pseudo-code of FedNGDB in Algorithm 2.
Appendix C Auxiliary Experiments
C.1. More Clients
@@ 3  Upload $\mathbf{E}^{i}$ to server for secret aggregation, receive $\mathbf{E}_{0}^{i}$
Table 6. The performance (MRR) of GQE when #C=10.
Setting FB15k-237 FB15k NELL995
Local 8.46 13.04 7.83
FedNGDB 10.17 14.98 9.17
C.2. Relation Overlap
@@
Table 7. The MRR of GQE when relation overlapped.
Setting FB15k-237 FB15k NELL995
Local 10.22 20.21 9.64
FedNGDB 11.42 22.47 11.36

### v4

@@ 6 Conclusion
Acknowledgement
The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong,SAR, China.
@@ Appendix A Parameter Sharing
In this section, we provide an example of sharing secrets between clients. The parameters can be shared under the protection using various encryption methods, for example, the commonly used is Diffie–Hellman key exchange (Diffie & Hellman, 2022) shown as follows, we consider the sharing process between two clients:
@@ 3  Upload $\mathbf{E}^{i}$ to server for secret aggregation, receive $\mathbf{E}_{0}^{i}$
B.2 FedNGDB Framework
We present the pseudo-code of FedNGDB in Algorithm 2.
B.3 Differential Privacy Setting
Differential privacy represents a mathematical framework for quantifying and limiting the disclosure of private information in statistical databases (Dwork, 2008). In federated learning, differential privacy enables collaborative model training across decentralized data sources while providing formal privacy guarantees. By injecting calibrated noise into the training process, these techniques ensure that individual contributions cannot be reliably identified or extracted from the global model (Geyer et al., 2017). The privacy-utility tradeoff can be tuned via the privacy budget parameter $\epsilon$, allowing practitioners to balance learning performance against disclosure risk. The formal definition of $\epsilon$-differential privacy ($\epsilon$-DP) is as follows (Dwork, 2006): A randomized algorithm $M$ satisfies $\epsilon$-differential privacy if for all neighboring datasets $D_{1}$ and $D_{2}$ that differ in at most one record, and for all possible outputs $S\subseteq Range(M)$:
$Pr[M(D_{1})\in S]\leq e^{\epsilon}Pr[M(D_{2})\in S]$
Where $\epsilon$ is the privacy parameter that quantifies the privacy loss.
In the former part, we introduce Laplacian noise to the local gradients to meet the guarantee of $\epsilon$-DP for in-graph query in local neural graph databases. However, for cross-graph query, we compute the score on each graph database for each query. The Post-Processing Theorem (Dwork et al., 2014) which states the composition of a data-independent mapping with an $\epsilon$-DP algorithm $M$ is also $\epsilon$-DP. Since the ranking operation of FedNGDB is a post-processing step applied to the outputs $(M_{1}(D),\cdots,M_{n}(D))$ from each local databases and the overlapped candidate answers share same scores, by the post-processing theorem of differential privacy, the final ranked list maintains the same privacy guarantee of $\epsilon$-DP.
Appendix C Auxiliary Experiments
C.1 More Clients
Table 5: The performance of GQE when #C=10.
FB15k-237 FB15k NELL995
In-graph Cross-graph In-graph Cross-graph In-graph Cross-graph
Graph Setting HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR
Local 8.96 8.47 - - 13.46 13.04 - - 8.36 7.83 - -
Central 13.15 12.76 13.24 12.89 19.76 19.43 20.27 19.87 13.76 13.28 13.53 13.27
GQE FedNGDB 10.67 10.17 10.23 10.04 15.39 14.98 14.86 14.45 12.93 12.11 12.55 11.96
C.2 Relation Overlap
@@
Table 6: The MRR of GQE when relation overlapped.
FB15k-237 FB15k NELL995
In-graph Cross-graph In-graph Cross-graph In-graph Cross-graph
Graph Setting HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR
Local 10.48 10.22 - - 20.76 20.21 - - 10.03 9.64 - -
Central 13.47 13.27 13.86 13.54 30.74 30.46 30.27 30.16 13.82 13.68 13.56 13.12
GQE FedNGDB 11.96 11.42 11.67 11.43 23.89 22.47 23.22 22.89 11.68 11.36 12.03 11.89
C.3 Convergence Rate
We evaluate the convergence speed of three federated frameworks. The results are presented by the average number of communication round ratios relative to FedE. As shown in Table 7, FedNGDB’s convergence speed is faster than FedR while slightly slower than FedE, demonstrating that our FedNGDB can protect stronger protection while remaining competitive efficiency.
Table 7: The statistics of communication rounds .
Setting FedE FedR FedNGDB
Relative Rounds to FedE 1.00 1.32 1.09